OpenStudy (anonymous):

Why is the following assertion not necessarily true: lim x->a f(x) = lim 2x->2a f(x)?

OpenStudy (anonymous):

A counterexample will suffice.

OpenStudy (jamesj):

Actually, this is true and I can prove it with epsilon-delta formality, if you are a college student who has seen it.

OpenStudy (anonymous):

I am a college student, and I have seen epsilon-delta proofs. Here's the official problem in my book: Give an example of a function f for which the following assertion is false: If |f(x)-l|<E when 0<|x-a|<d, then |f(x)-l<E/2 when 0<|x-a|<d/2. (Where E is epsilon and d is delta). The first half seems to imply that lim x->a f(x)=l, and the second half (which I'm supposed to show is false) seems to imply that lim 2x->2a 2f(x)=2l; we factor out a 2 from each side, and we arrive at lim 2x->2a f(x)=l. Am I doing anything wrong?

OpenStudy (jamesj):

Ah ... ok. The problem from your book is something else and it is true.

OpenStudy (anonymous):

Oh... what did I do wrong?

OpenStudy (jamesj):

by which I mean it is indeed the case that if .... < d => ... < e then you can't necessarily say that ... < d/2 => .... < e/2

OpenStudy (anonymous):

Ah, okay. Any function you can think of that makes the statement false?

OpenStudy (jamesj):

For example f(x) = x^(1/3) . Look at the limit as x --> 0

OpenStudy (jamesj):

Now it is indeed true that for all e > 0 there exists a d such that |x-0| < d => |x^(1/3) - 0|< e You can see how intuitively this is going to fail because for x close to zero, the function approaches 0 slowly. So if you halved the distance to zero, i.e., looked at d/2, it would not necessarily be the case that |x^(1/3)| < e

OpenStudy (jamesj):

To formalize this idea, start as always with an e, let it be 0 < e < 1. Then find explicitly the tightest value of d you can for which |x-0| < d => |x^(1/3) - 0|< e Then show it is false that |x-0| < d/2 => |x^(1/3) - 0|< e/2

OpenStudy (jamesj):

Yes?

OpenStudy (anonymous):

Hm... one sec...

OpenStudy (anonymous):

Oh, okay. Let me see...

OpenStudy (anonymous):

I still don't really see how that works. What do you mean with the "tightest" value of d? The smallest value of d, so that we restrict x as much as possible?

OpenStudy (jamesj):

Actually let's got back to f(x) = x^(1/3) For e > 0 and e < 1, then for d = e^3 |x - 0| < d => |x^(1/3) - 0| < 0

OpenStudy (anonymous):

< 0?

OpenStudy (anonymous):

Where did that come from?

OpenStudy (jamesj):

< e But if |x - 0|<d/2 = e^3/2, then it is NOT true that |x^(1/3)|<e/2 because |x|<e^3 / 2 => |x^(1/3)| < e/2^(1/3) and 1/2^(1/3) > 1/2

OpenStudy (jamesj):

There's some formalism here, but do you understand the idea? Draw the graph of f(x) = x^(1/3) over the domain [-1,1] and see how this function is behaving.

OpenStudy (anonymous):

That all makes sense except for the last statement, I have no idea where that came from.

OpenStudy (jamesj):

Just literally taking the cube root of both sides of |x|<e^3 / 2

OpenStudy (anonymous):

Then we'd get x^(1/3)<e/2^(1/3)? I still don't see what you did

OpenStudy (jamesj):

The point is \[\frac{\epsilon}{2^{1/3}} > \frac{\epsilon}{2}\]

OpenStudy (jamesj):

We can't get the sharpness or tightness of the inequality that we need. To make it explicity, choose an epsilon. Let's take epsilon, e= 1/1000

OpenStudy (jamesj):

then for d= 1/10 |x - 0| < d => |x^(1/3) - 0| < e

OpenStudy (anonymous):

Sure.

OpenStudy (jamesj):

But for d/2 = 1/20. It is the case that ....

OpenStudy (anonymous):

1/8000 < e

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

Scuse me, 1/8000<e/2

OpenStudy (jamesj):

Oh, my first qualm was right. Sorry. Take f(x) = x^2 instead. And let's take epsilon = 1/100. Then for delta, d = 1/10 |x - 0| < d => |x^2 - 0| < e

OpenStudy (jamesj):

No. Stay with our first example: f(x) = x^1/3

OpenStudy (anonymous):

Haha, ok.

OpenStudy (jamesj):

So for f(x) = x^(1/3). Take e = 1/10. Then for d = 1/1000 |x-0| < d => |f(x)| < e

OpenStudy (jamesj):

That's what got us caught up above

OpenStudy (jamesj):

Now, for d/2 = 1/2000 |x| < d/2 => |f(x)| < 1/2000^(1/3) = 1/(2^(1/3).10)

OpenStudy (anonymous):

What does 1/(2^(1/3).10) mean?

OpenStudy (anonymous):

Never mind actually, I get that.

OpenStudy (jamesj):

cube root of 2 times 10. Now to show it is not true that |x| < d/2 => |f(x)|<e/2 we need to show that there is an x such taht |x| < d/2 and |f(x))| >= e/2

OpenStudy (jamesj):

so we're looking for an x such that |x| < 1/2000 and x^(1/3) >= 1/20

OpenStudy (jamesj):

and of course there infinity many such x because that last inequality is equivalent to x >= 1/8000

OpenStudy (anonymous):

Ah.

OpenStudy (anonymous):

You have no idea how confused I am right now but I'm going to go read over this and see if I can make sense out of it. Thanks.

OpenStudy (jamesj):

Like I said above somewhere, draw the graph of f(x) = x^(1/3) and sketch out the delta/epsilon inequalities on the graph and you'll see why this works.

OpenStudy (anonymous):

All right, that makes sense. But how did you know to take f(x)=x^(1/3) and what values to pick for e and d?

OpenStudy (jamesj):

You can see that this would work for a linear function like f(x) = ax

OpenStudy (jamesj):

So we needed a function that approaches something slower than linearly. Again, the graph gives you the picture and if we were in a room with a white board--vs. the blah drawing tool here--I'd show you.

OpenStudy (jamesj):

Anyway ... sorry for a bit a confusion there.

OpenStudy (anonymous):

Ahhh, ok. Thanks a lot.